Bad weather kept us closed until about 02:00 MDT. During preflight checks, we had a recurrence of the secondary E actuator failing to finish its move on a full step, so that the collimate process reported errors. We homed the E actuator (the position error was 1), which solved the problem. We moved the mirror arround at various altitudes and could not reproduce the error. The PR (2560) was updated. We opened under partly cloudy skies at ~02:00 and tested the pointing model installed during shakedown. We focx We observed a 20-point grid (TGRID:GRID_5_4_25_80_T_30.DAT) then used tpoint to calcuate the mean error between the observed and fitted positions.